# Integral representation of solutions to initial-boundary value problems in the framework of the hyperbolic heat equation

### v1

@@ 3.4 Deformation of the integration path
In order to obtain a genuine representation of the solution, we first need to deform the integration path in the integrals involving the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ in Eq. (3.12) from the real axis to appropriate paths in the complex $k$-plane. We define the domains $D^{\pm}=D\cap\mathbb{C}^{\pm}$, where ±=$\mathbb{C}^{\pm}=\{k\in\mathbb{C}:\operatorname{Im}k\gtrless 0\}$, and
$D=\{k\in\mathbb{C}:\operatorname{Re}\omega_{1}(k)<0\text{ or }\operatorname{Re}\omega_{2}(k)<0\}\equiv\left\{k=k_{\mathrm{R}}+\mathrm{i}k_{\mathrm{I}}:k_{\mathrm{I}}^{2}\geq\frac{k_{\mathrm{R}}^{2}}{4\alpha\tau k_{\mathrm{R}}^{2}+1}\right\}.$
@@
where
$\partial D^{\pm}=\left\{k=k_{\mathrm{R}}+\mathrm{i}k_{\mathrm{I}}:k_{\mathrm{I}}=\pm\sqrt{\frac{k_{\mathrm{R}}^{2}}{4\alpha\tau k_{\mathrm{R}}^{2}+1}},\ k_{\mathrm{R}}\in(-\infty,\infty)\right\},$
and we assume that the path of integration along the boundary of a domain is directed in such a way that the domain is to the left of the boundary when it is traversed. Besides, we deform the paths of integration from $\partial D^{\pm}$ to $\partial D_{0}^{\pm}$ in the vicinity of the origin. Note that the imaginary part of the spectral parameter $k$ on the integration paths is bounded.
@@ 3.5 Using the symmetry of the dispersion relations
$\gamma_{0}\tilde{e}^{(m)}(0,t)+\delta_{0}\tilde{q}^{(m)}(0,t)=\tilde{g}^{(m)}(t),\quad\gamma_{l}\tilde{e}^{(m)}(l,t)+\delta_{l}\tilde{q}^{(m)}(l,t)=\tilde{h}^{(m)}(t),$ (3.18)
@@
To avoid too cumbersome general calculations, below we will limit ourselves to the most interesting boundary conditions from a physical point of view. We will limit ourselves to boundary conditions that determine the heat flux or express Newton’s law. We will not consider the conditions determining the boundary temperature. Therefore, we assume that $\delta_{0}=1$, $\delta_{l}=-1$ and $\gamma_{0,l}\geq 0$. In this case the boundary conditions (2.3) take the form
@@
$\tilde{q}^{(m)}(0,t)=\tilde{g}^{(m)}(t)-\gamma_{0}\tilde{e}^{(m)}(0,t),\quad\tilde{q}^{(m)}(l,t)=\gamma_{l}\tilde{e}^{(m)}(l,t)-\tilde{h}^{(m)}(t).$ (3.21)
Using the relations (3.21) we can express the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ (Eq. (3.8)) via $\tilde{e}^{(m)}$, $\tilde{g}^{(m)}$ and $\tilde{h}^{(m)}$:
$\boldsymbol{a}(k,t)=\begin{pmatrix}-\sigma_{0}^{(2)}(k)\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{g}^{(1)}(t)\\[4.0pt] \sigma_{0}^{(1)}(k)\tilde{e}^{(2)}(0,t)-\omega_{1}\tilde{g}^{(2)}(t)\end{pmatrix},$ (3.22a)
$\boldsymbol{b}(k,t)=\begin{pmatrix}\sigma_{l}^{(2)}(-k)\tilde{e}^{(1)}(l,t)-\omega_{2}\tilde{h}^{(1)}(t)\\[4.0pt] -\sigma_{l}^{(1)}(-k)\tilde{e}^{(2)}(l,t)+\omega_{1}\tilde{h}^{(2)}(t)\end{pmatrix},$ (3.22b)
@@
$\sigma_{0,l}^{(m)}(k)=\mathrm{i}k\beta+\gamma_{0,l}\omega_{m}(k).$ (3.23)
The vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ depend on the functions $\tilde{e}^{(1,2)}$, which are still unknown. To find them, we use the system (3.17) of the global relations and the representations (3.22). The system (3.17), consisting of four equations, can be rewritten as the system of two algebraic systems
$\begin{pmatrix}\sigma_{0}^{(2)}(k)&\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[4.0pt] \sigma_{0}^{(2)}(-k)&\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\end{pmatrix}\begin{pmatrix}\tilde{e}^{(1)}(0,t)\\[4.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}$ $=\begin{pmatrix}U_{1}(k,t)-V_{1}(k,t)\\[4.0pt] U_{1}(-k,t)-V_{1}(-k,t)\end{pmatrix},$ (3.24a)
$\begin{pmatrix}\sigma_{0}^{(1)}(k)&\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[2.0pt] \sigma_{0}^{(1)}(-k)&\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\end{pmatrix}\begin{pmatrix}\tilde{e}^{(2)}(0,t)\\[2.0pt] \tilde{e}^{(2)}(l,t)\end{pmatrix}$ $=-\begin{pmatrix}U_{2}(k,t)-V_{2}(k,t)\\[2.0pt] U_{2}(-k,t)-V_{2}(-k,t)\end{pmatrix}$ (3.24b)
@@
$\begin{split}V_{1}(k,t)=\varPhi_{1}(k)+\tilde{F}_{1}(k,t)-\omega_{2}\left[\tilde{g}^{(1)}(t)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\tilde{h}^{(1)}(t)\right],\\ V_{2}(k,t)=\varPhi_{2}(k)+\tilde{F}_{2}(k,t)+\omega_{1}\left[\tilde{g}^{(2)}(t)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\tilde{h}^{(2)}(t)\right],\end{split}$ (3.25)
$\varPhi_{1,2}$ are given by Eq. (3.16), $\tilde{F}_{1,2}$ are given by Eq. (3.11), $\tilde{g}^{(1,2)}$ and $\tilde{h}^{(1,2)}$ are given by Eq. (3.19). The inverse of the matrices in the systems (3.24a) and (3.24b) are given by
@@
Solving the systems (3.24), we find the functions $\tilde{e}^{(1,2)}$:
$\begin{pmatrix}\tilde{e}^{(1)}(0,t)\\[2.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}=\frac{1}{\varDelta_{1}(k)}\Bigg{[}$ $\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\\[2.0pt] -\sigma_{0}^{(2)}(-k)\end{pmatrix}\left[U_{1}(k,t)-V_{1}(k,t)\right]$ (3.27a)
$+\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[2.0pt] \sigma_{0}^{(2)}(k)\end{pmatrix}\left[U_{1}(-k,t)-V_{1}(-k,t)\right]\Bigg{]},$
$\begin{pmatrix}\tilde{e}^{(2)}(0,t)\\[2.0pt] \tilde{e}^{(2)}(l,t)\end{pmatrix}=\frac{1}{\varDelta_{2}(k)}\Bigg{[}$ $\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\\[2.0pt] \sigma_{0}^{(1)}(-k)\end{pmatrix}\left[U_{2}(k,t)-V_{2}(k,t)\right]$ (3.27b)
$+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[2.0pt] -\sigma_{0}^{(1)}(k)\end{pmatrix}\left[U_{2}(-k,t)-V_{2}(-k,t)\right]\Bigg{]},$
where $U_{1,2}$ are given by Eq. (3.15) and $V_{1,2}$ are given by Eq. (3.25). Note that the functions $U_{1,2}$ are still unknown, while the functions $V_{1,2}$ are known, they contain only known functions.
@@ 3.6 The solution
The formula (3.13) is also not the final representation of the solution, because the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ depend on the unknown functions $e$ and $q$ through the functions $U_{1,2}$ (see Eqs. (3.22), (3.27) and (3.15)). However, it can be shown that integrals containing the functions $U_{1,2}$ are equal to zero, see, e. g., Refs. [25, 27, 29, 30], and, therefore, the final formula does not contain these integrals.
@@
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)(t-s)}\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}k(l-x)}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (3.28)
where $\hat{\boldsymbol{u}}_{0}(k)$ is given by Eq. (3.7), $\hat{\boldsymbol{F}}(k,t)$ is given by Eq. (3.10),
$\boldsymbol{a}^{\prime}(k,t)=\begin{pmatrix}-\sigma_{0}^{(2)}(k)\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{g}^{(1)}(t)\\[4.0pt] \sigma_{0}^{(1)}(k)\tilde{e}^{(2)}(0,t)-\omega_{1}\tilde{g}^{(2)}(t)\end{pmatrix},$ (3.29a)
$\boldsymbol{b}^{\prime}(k,t)=\begin{pmatrix}\sigma_{l}^{(2)}(-k)\tilde{e}^{(1)}(l,t)-\omega_{2}\tilde{h}^{(1)}(t)\\[4.0pt] -\sigma_{l}^{(1)}(-k)\tilde{e}^{(2)}(l,t)+\omega_{1}\tilde{h}^{(2)}(t)\end{pmatrix},$ (3.29b)
(the expressions for $\boldsymbol{a}^{\prime}$ and $\boldsymbol{b}^{\prime}$ formally coincide with the expressions for $\boldsymbol{a}$ and $\boldsymbol{b}$, see Eq. (3.22), however, the functions $\tilde{e}^{(1,2)}$ are different, see Eq. (3.30) below), the coefficients $\sigma_{0,l}^{(1,2)}$ are given by Eq. (3.23), the functions $\tilde{g}^{(1,2)}$ and $\tilde{h}^{(1,2)}$ are given by Eq. (3.19), the functions $\tilde{e}^{(1,2)}$ are given by
$\begin{pmatrix}\tilde{e}^{(1)}(0,t)\\[3.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{1}(k)}\Bigg{[}\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\\[3.0pt] \sigma_{0}^{(2)}(-k)\end{pmatrix}V_{1}(k,t)+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[3.0pt] -\sigma_{0}^{(2)}(k)\end{pmatrix}V_{1}(-k,t)\Bigg{]},$ (3.30a)
$\begin{pmatrix}\tilde{e}^{(2)}(0,t)\\[3.0pt] \tilde{e}^{(2)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{2}(k)}\Bigg{[}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\\[3.0pt] -\sigma_{0}^{(1)}(-k)\end{pmatrix}V_{2}(k,t)+\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[3.0pt] \sigma_{0}^{(1)}(k)\end{pmatrix}V_{2}(-k,t)\Bigg{]}$ (3.30b)
(these are Eqs. (3.27) with $U_{1,2}=0$), the functions $V_{1,2}$ are given by Eq. (3.25), the coefficients $\varDelta_{1,2}(k)$ are given by Eq. (3.26).

### v2

@@ 3.4 Deformation of the integration path
In order to obtain a genuine representation of the solution, we first need to deform the integration path in the integrals involving the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ in Eq. (3.12) from the real axis to appropriate paths in the complex $k$-plane. We define the domains $D^{\pm}=D\cap\mathbb{C}^{\pm}$, where $\mathbb{C}^{\pm}=\{k\in\mathbb{C}:\operatorname{Im}k\gtrless 0\}$, and
$D=\{k\in\mathbb{C}:\operatorname{Re}\omega_{1}(k)<0\text{ or }\operatorname{Re}\omega_{2}(k)<0\}.$
@@
where the boundaries $\partial D^{\pm}$ of the domains $D^{\pm}$ are given by
$\partial D^{\pm}=\left\{k=k_{\mathrm{R}}+\mathrm{i}k_{\mathrm{I}}:k_{\mathrm{I}}=\pm\sqrt{\dfrac{\big(\mu^{2}|k|^{2}-1\big)^{2}+4\mu^{2}k_{\mathrm{R}}^{2}}{\big(\mu^{2}|k|^{2}-1\big)^{2}+4C\mu^{2}k_{\mathrm{R}}^{2}}}|k_{\mathrm{R}}|,\ k_{\mathrm{R}}\in(-\infty,\infty)\right\},$ (3.14)
$C=\alpha\tau/\mu^{2}$ (see Appendix A), and we assume that the path of integration along the boundary of a domain is directed in such a way that the domain is to the left of the boundary when it is traversed. Besides, we deform the paths of integration from $\partial D^{\pm}$ to $\partial D_{0}^{\pm}$ in the vicinity of the origin. Note that if $C=1$, then the boundaries are given by
$\partial D^{\pm}=\left\{k=k_{\mathrm{R}}+\mathrm{i}k_{\mathrm{I}}:k_{\mathrm{I}}=\pm|k_{\mathrm{R}}|,\ k_{\mathrm{R}}\in(-\infty,\infty)\right\}.$
Eq. (3.14) implies also that
$\left|\frac{k_{\mathrm{I}}}{k_{\mathrm{R}}}\right|\to 1\quad\text{as}\quad k_{\mathrm{R}}\to 0\enskip\text{or}\enskip k_{\mathrm{R}}\to\infty,$
see Appendix A.
@@ 3.5 Using the symmetry of the dispersion relations
$\gamma_{0}\tilde{e}_{0}^{(m)}(0,t)+\delta_{0}\tilde{q}_{0}^{(m)}(0,t)=\tilde{g}^{(m)}(t),\quad\gamma_{l}\tilde{e}_{0}^{(m)}(l,t)+\delta_{l}\tilde{q}_{0}^{(m)}(l,t)=\tilde{h}^{(m)}(t),$ (3.19)
@@
To avoid too cumbersome general calculations, below we will limit ourselves to the most interesting boundary conditions from a physical point of view. We will limit ourselves to boundary conditions that determine the heat flux or express Newton’s law. We will not consider the conditions determining the boundary temperature. Therefore, we assume that $\delta_{0}=1$, $\delta_{l}=-1$ and $\gamma_{0,l}\geq 0$. In this case the boundary conditions take the form
@@
$\tilde{q}_{0}^{(m)}(0,t)=\tilde{g}^{(m)}(t)-\gamma_{0}\tilde{e}_{0}^{(m)}(0,t),\quad\tilde{q}_{0}^{(m)}(l,t)=\gamma_{l}\tilde{e}_{0}^{(m)}(l,t)-\tilde{h}^{(m)}(t).$ (3.22)
Besides, Eqs. (2.1a) and (2.2) imply the relations
$\tilde{q}_{1}^{(m)}(x,t)=\omega_{m}\tilde{e}_{0}^{(m)}(x,t)-e(x,t)+\varphi(x)\mathop{\mathrm{e}}\nolimits^{-\omega_{m}t}+\tilde{f}^{(m)}(x,t),$ (3.23)
@@
$\tilde{f}^{(m)}(x,t)\equiv\tilde{f}(\omega_{m}(k),x,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}f(x,s)\mathop{}\!\mathrm{d}s$ (3.24)
is the time transform of $f$.
Using the relations (3.22) and (3.23) we can express the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ (Eq. (3.8)) via $\tilde{e}_{0}^{(1,2)}$, $\tilde{g}^{(m)}$ and $\tilde{h}^{(m)}$:
$\boldsymbol{a}(k,t)=$ $\begin{pmatrix}-\sigma_{0}^{(1)}(k)\tilde{e}_{0}^{(1)}(0,t)+(\omega_{2}-\theta k^{2})\tilde{g}^{(1)}(t)\\[3.0pt] \sigma_{0}^{(2)}(k)\tilde{e}_{0}^{(2)}(0,t)-(\omega_{1}-\theta k^{2})\tilde{g}^{(2)}(t)\end{pmatrix}+\mathrm{i}k\theta\begin{pmatrix}-e(0,t)+L_{1}(k,t)\\[3.0pt] e(0,t)-L_{2}(k,t)\end{pmatrix}$, (3.25a)
$\boldsymbol{b}(k,t)=$ $\begin{pmatrix}\sigma_{l}^{(1)}(-k)\tilde{e}_{0}^{(1)}(l,t)-(\omega_{2}-\theta k^{2})\tilde{h}^{(1)}(t)\\[3.0pt] -\sigma_{l}^{(2)}(-k)\tilde{e}_{0}^{(2)}(l,t)+(\omega_{1}-\theta k^{2})\tilde{h}^{(2)}(t)\end{pmatrix}+\mathrm{i}k\theta\begin{pmatrix}-e(l,t)+R_{1}(k,t)\\[3.0pt] e(l,t)-R_{2}(k,t)\end{pmatrix}$, (3.25b)
where
$\begin{split}\sigma_{0,l}^{(1)}(k)=\mathrm{i}k(\beta-\theta\omega_{1})+\gamma_{0,l}(\omega_{2}-\theta k^{2}),\\ \sigma_{0,l}^{(2)}(k)=\mathrm{i}k(\beta-\theta\omega_{2})+\gamma_{0,l}(\omega_{1}-\theta k^{2}),\end{split}$ (3.26)
$\begin{split}L_{1}(k,t)=\varphi(0)\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}+\tilde{f}^{(1)}(0,t),\quad R_{1}(k,t)=\varphi(l)\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}+\tilde{f}^{(1)}(l,t),\\ L_{2}(k,t)=\varphi(0)\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t}+\tilde{f}^{(2)}(0,t),\quad R_{2}(k,t)=\varphi(l)\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t}+\tilde{f}^{(2)}(l,t).\end{split}$ (3.27)
The vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ depend on the functions $\tilde{e}_{0}^{(1,2)}$, which are still unknown. To find them, we use the system (3.18) of the global relations and the representations (3.25). The system (3.18), consisting of four equations, can be rewritten as the system of two algebraic systems
$\begin{pmatrix}\sigma_{0}^{(1)}(k)&\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[4.0pt] \sigma_{0}^{(1)}(-k)&\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\end{pmatrix}\begin{pmatrix}\tilde{e}_{0}^{(1)}(0,t)\\[4.0pt] \tilde{e}_{0}^{(1)}(l,t)\end{pmatrix}$ (3.28a)
$\hskip 90.0pt=\begin{pmatrix}U_{1}(k,t)-V_{1}(k,t)-W_{1}(k,t)-E(k,t)\\[4.0pt] U_{1}(-k,t)-V_{1}(-k,t)-W_{1}(-k,t)-E(-k,t)\end{pmatrix}$,
$\begin{pmatrix}\sigma_{0}^{(2)}(k)&\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[2.0pt] \sigma_{0}^{(2)}(-k)&\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\end{pmatrix}\begin{pmatrix}\tilde{e}_{0}^{(2)}(0,t)\\[2.0pt] \tilde{e}_{0}^{(2)}(l,t)\end{pmatrix}$ (3.28b)
$\hskip 90.0pt=-\begin{pmatrix}U_{2}(k,t)-V_{2}(k,t)-W_{2}(k,t)-E(k,t)\\[2.0pt] U_{2}(-k,t)-V_{2}(-k,t)-W_{2}(-k,t)-E(-k,t)\end{pmatrix}$
@@
$\begin{split}V_{1}(k,t)&=\varPhi_{1}(k)+\tilde{F}_{1}(k,t)-(\omega_{2}-\theta k^{2})\big[\tilde{g}^{(1)}(t)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\tilde{h}^{(1)}(t)\big],\\ V_{2}(k,t)&=\varPhi_{2}(k)+\tilde{F}_{2}(k,t)+(\omega_{1}-\theta k^{2})\big[\tilde{g}^{(2)}(t)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\tilde{h}^{(2)}(t)\big],\end{split}$ (3.29)
$\varPhi_{1,2}$ are given by Eq. (3.17), $\tilde{F}_{1,2}$ are given by Eq. (3.10), $\tilde{g}^{(1,2)}$ and $\tilde{h}^{(1,2)}$ are given by Eq. (3.20),
$\begin{split}W_{1}(k,t)&=-\mathrm{i}k\theta\left[L_{1}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}R_{1}(k,t)\right]\\ &\equiv-\mathrm{i}k\theta\left[\varphi(0)\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}+\tilde{f}^{(1)}(0,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\left(\varphi(l)\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}+\tilde{f}^{(1)}(l,t)\right)\right],\\ W_{2}(k,t)&=\mathrm{i}k\theta\left[L_{2}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}R_{2}(k,t)\right]\\ &\equiv\mathrm{i}k\theta\left[\varphi(0)\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t}+\tilde{f}^{(2)}(0,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\left(\varphi(l)\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t}+\tilde{f}^{(2)}(l,t)\right)\right],\end{split}$
and
$E(k,t)=\mathrm{i}k\theta\big[e(0,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}e(l,t)\big].$ (3.30)
The inverse of the matrices in the systems (3.28a) and (3.28b) are given by
@@
Solving the systems (3.28), we find the functions $\tilde{e}_{0}^{(1,2)}$:
$\begin{pmatrix}\tilde{e}_{0}^{(1)}(0,t)\\[2.0pt] \tilde{e}_{0}^{(1)}(l,t)\end{pmatrix}$= $\frac{1}{\varDelta_{1}(k)}\Bigg[\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\\[2.0pt] -\sigma_{0}^{(1)}(-k)\end{pmatrix}\left[U_{1}(k,t)-V_{1}(k,t)-W_{1}(k,t)-E(k,t)\right]$ (3.32a)
$+\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[2.0pt] \sigma_{0}^{(1)}(k)\end{pmatrix}\left[U_{1}(-k,t)-V_{1}(-k,t)-W_{1}(-k,t)-E(-k,t)\right]\Bigg],$
$\begin{pmatrix}\tilde{e}_{0}^{(2)}(0,t)\\[2.0pt] \tilde{e}_{0}^{(2)}(l,t)\end{pmatrix}$= $\frac{1}{\varDelta_{2}(k)}\Bigg[\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\\[2.0pt] \sigma_{0}^{(2)}(-k)\end{pmatrix}\left[U_{2}(k,t)-V_{2}(k,t)-W_{2}(k,t)-E(k,t)\right]$ (3.32b)
$+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[2.0pt] -\sigma_{0}^{(2)}(k)\end{pmatrix}\left[U_{2}(-k,t)-V_{2}(-k,t)-W_{2}(-k,t)-E(-k,t)\right]\Bigg],$
where $U_{1,2}$ are given by Eq. (3.16), $V_{1,2}$ are given by Eq. (3.29), $W_{1,2}$ are given by Eq. (3.5) and $E$ is given by Eq. (3.30). Note that the functions $U_{1,2}$ and $E$ are still unknown, while the functions $V_{1,2}$ and $W_{1,2}$ are known, they contain only known functions.
@@ 3.6 The solution
The formula (3.13) is also not the final representation of the solution, because the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ depend on the unknown functions $e$ and $q$ through the functions $U_{1,2}$ and $E$ (see Eqs. (3.25), (3.32), (3.16) and (3.30)). However, it can be shown that integrals containing the functions $\hat{e}(k,t)$, $\hat{q}(k,t)$, $e(0,t)$ and $e(l,t)$ are equal to zero, see, e. g., Refs. [20, 22, 24, 25], and, therefore, the final formula does not contain these integrals.
@@
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)(t-s)}\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (3.33)
where $\hat{\boldsymbol{u}}_{0}(k)$ is given by Eq. (3.7), $\hat{\boldsymbol{F}}(k,t)$ is given by Eq. (3.11), the vectors $\boldsymbol{a}^{\prime}$ and $\boldsymbol{b}^{\prime}$ are given by
$\boldsymbol{a}^{\prime}(k,t)=\begin{pmatrix}-\sigma_{0}^{(1)}(k)\tilde{e}_{0}^{(1)}(0,t)+(\omega_{2}-\theta k^{2})\tilde{g}^{(1)}(t)\\[3.0pt] \sigma_{0}^{(2)}(k)\tilde{e}_{0}^{(2)}(0,t)-(\omega_{1}-\theta k^{2})\tilde{g}^{(2)}(t)\end{pmatrix}+\mathrm{i}k\theta\begin{pmatrix}L_{1}(k,t)\\[3.0pt] -L_{2}(k,t)\end{pmatrix}$, (3.34a)
$\boldsymbol{b}^{\prime}(k,t)=\begin{pmatrix}\sigma_{l}^{(1)}(-k)\tilde{e}_{0}^{(1)}(l,t)-(\omega_{2}-\theta k^{2})\tilde{h}^{(1)}(t)\\[3.0pt] -\sigma_{l}^{(2)}(-k)\tilde{e}_{0}^{(2)}(l,t)+(\omega_{1}-\theta k^{2})\tilde{h}^{(2)}(t)\end{pmatrix}+\mathrm{i}k\theta\begin{pmatrix}R_{1}(k,t)\\[3.0pt] -R_{2}(k,t)\end{pmatrix}$ (3.34b)
(these are Eqs. (3.25) with $e(0,t)=0$ and $e(l,t)=0$ ), the functions $\tilde{e}_{0}^{(1,2)}$ are given by
$\begin{pmatrix}\tilde{e}_{0}^{(1)}(0,t)\\[2.0pt] \tilde{e}_{0}^{(1)}(l,t)\end{pmatrix}=\frac{1}{\varDelta_{1}(k)}\Bigg[$ $\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\\[2.0pt] \sigma_{0}^{(1)}(-k)\end{pmatrix}\left[V_{1}(k,t)+W_{1}(k,t)\right]$ (3.35a)
$+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[2.0pt] -\sigma_{0}^{(1)}(k)\end{pmatrix}\left[V_{1}(-k,t)+W_{1}(-k,t)\right]\Bigg],$
$\begin{pmatrix}\tilde{e}_{0}^{(2)}(0,t)\\[2.0pt] \tilde{e}_{0}^{(2)}(l,t)\end{pmatrix}=\frac{1}{\varDelta_{2}(k)}\Bigg[$ $\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\\[2.0pt] -\sigma_{0}^{(2)}(-k)\end{pmatrix}\left[V_{2}(k,t)+W_{2}(k,t)\right]$ (3.35b)
$+\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[2.0pt] \sigma_{0}^{(2)}(k)\end{pmatrix}\left[V_{2}(-k,t)+W_{2}(-k,t)\right]\Bigg]$
(these are Eqs. (3.32) with $U_{1,2}=0$ and $E=0$).